A conducting disc of radius R rotates about its axis with an Angular velocity. Then the potential difference between the centre of the disc and its edge is (no magnetic field is present) :

Grade:11

2 Answers

Gitanjali Rout
5 years ago
a)Free electrons of metal disc are rotating, so there has to be some force that is responsible for centripetal acceleration of electrons. That force exists because free electrons go to the rim of the disc and create electric field, so we have: Fcp=e*E, E=m*ω^2*r/e. So potential diference is Δφ=∫Edr (from 0 to R), Δφ=mω^2*r^2/2e.

Reference
thank you
 
Gitanjali Rout
5 years ago
a)Free electrons of metal disc are rotating, so there has to be some force that is responsible for centripetal acceleration of electrons. That force exists because free electrons go to the rim of the disc and create electric field, so we have: Fcp=e*E, E=m*ω^2*r/e. So potential diference is Δφ=∫Edr (from 0 to R), Δφ=mω^2*r^2/2e.